OpenStudy (anonymous):

2(y + 7) = 4y + 14 is it y = 0??

OpenStudy (nenadmatematika):

yes :D

hero (hero):

2(y+7) = 4y + 14 2y+ 14 = 4y+14 2y = 4y 0 = 4y - 2y 0 = 2y 0 = y

OpenStudy (nenadmatematika):

Chinagirl has troubles with the given choices of answers....It seems that the correct solution is not among the given choices...:D chinagirl have faith in yourself :D